OpenStudy (anonymous):

State F.O.I.L methood using your actual example and show how you would use it to find the standard formula.

OpenStudy (anonymous):

first outside inside last

OpenStudy (anonymous):

this could be used in a problem such as 10x+12x+6

OpenStudy (anonymous):

Thank you marlins0412! :) Saved my life!

OpenStudy (anonymous):

your welcome what did you need this for?

OpenStudy (anonymous):

Working on a math worksheet for school. Teacher taught this to us today.

OpenStudy (anonymous):

This month* not today

OpenStudy (anonymous):

oh ok cool im guessing 10th grade or 9th?

OpenStudy (anonymous):

8th

OpenStudy (anonymous):

wow im in 10th and this was the first chapter in the book.

OpenStudy (anonymous):

Haha, my school is advanced in most subjects, i got algebra II next year i think

OpenStudy (anonymous):

wow thats really good.alright well off to my homework aswell se ya around!

OpenStudy (anonymous):

Good luck! Haha :)

OpenStudy (anonymous):

\[\begin{align} \text{F.O.I.L. = First Outside Inside Last}\\ \hline \\ &(ax+b)(cx+d)\\ \ \\ \text{First:}&(\color{blue}{ax}+b)(\color{blue}{cx}+d)\to \color{blue}{ax}\cdot\color{blue}{cx}=\color{blue}{acx^2}\\ \text{Outside:}&(\color{green}{ax}+b)(cx+\color{green}{d})\to \color{green}{ax}\cdot\color{green}{d}=\color{green}{adx}\\ \text{Inside:}&(ax+\color{red}{b})(\color{red}{cx}+d)\to \color{red}{b}\cdot\color{red}{cx}=\color{red}{bcx}\\ \text{Last:}&(ax+\color{orange}{b})(cx+\color{orange}{d})\to \color{orange}{b}\cdot\color{orange}{d}=\color{orange}{bd}\\ \hline \\ (ax+b)(cx+d)&=\color{blue}{acx^2}+\color{green}{adx}+\color{red}{bcx}+\color{orange}{bd} \end{align}\]